# Semiparametric Double Reinforcement Learning with Applications to Long-Term Causal Inference

### v3

@@ 4 Methods and challenges in nuisance estimation
4:     Update $q_{n,H}^{\pi,(k+1)}\in H$ by estimating:
@@
6:  Set $q_{n,H}:=q_{n,H}^{\pi,(K)}$;
7:  return  $q_{n,H}$;
One commonly used approach to estimate the Riesz representer $\alpha_{0,H}$ is to recast the minimization objective in (5) as the following convex-concave min-max optimization problem (Liu et al., 2018; Uehara et al., 2020; Dikkala et al., 2020; Kallus and Uehara, 2020):
$\alpha_{0,H}=\operatorname*{argmin}_{\alpha\in H}\max_{f\in L^{2}(P_{S_{0},A})}L_{0}(\alpha,f),$ (9)
where the objective function is given by:
$L_{0}(\alpha,f)=E_{0}\Big{[}\{\alpha(A_{0},S_{0})\}^{2}-$ $2\gamma\alpha(A_{0},S_{0})V^{\pi}(\alpha)(S_{1})-2m(S_{0},A_{0},\alpha)$
$-\frac{\gamma^{2}}{2}\left[\{f(A_{0},S_{0})\}^{2}-2V^{\pi}(\alpha)(S_{1})f(A_{0},S_{0})\right]\Big{]}.$
Given an estimator $\alpha_{n,H}$ of $\alpha_{0,H}$, the remaining nuisance component in $\mathcal{T}_{0}(\alpha_{0,H})$ can be estimated by regressing $V^{\pi}(\alpha_{n,H})(S_{1})$ on $(A_{0},S_{0})$, or equivalently, by computing $\arg\max_{f}L_{0}(\alpha_{n,H},f)$. Minimax estimation of solutions to conditional moment restrictions using empirical risk minimization techniques has been studied in Dikkala et al. (2020) and Bennett et al. (2023a).
To mitigate the computational challenges of min–max optimization, a common approach is to replace the inner unconstrained maximization with a constrained maximization over a linear model or a reproducing kernel Hilbert space (RKHS). In this case, the inner maximization can be efficiently computed in closed form using methods such as kernel ridge regression (Kallus and Uehara, 2020). The min–max optimization then reduces to a standard minimization problem based on the profiled loss (Murphy and Van der Vaart, 2000). A limitation of this approach is that constraining the inner maximization to a linear model or RKHS, while computationally attractive, can introduce substantial bias if the model is misspecified.
5 Automatic debiasing via Bellman Calibration
5.1 Leveraging the Q-function for dimension reduction
A central challenge in DRL is estimating the Riesz representer $\mathcal{T}_{0}(\alpha_{0,H})$ of the linear functional, such as the density ratio $d_{0}$ in (2). This quantity must typically be estimated by solving the min–max problem in (9), which is both computationally intensive and potentially unstable. To address this issue, we propose a specific instance of the adaptive DRL estimator developed in Section 3.4, which leverages a data-adaptive dimension reduction based on the $Q$-function. This approach only requires estimation of the $Q$-function, which can be readily obtained via fitted $Q$-iteration, and avoids estimating the Riesz representer altogether. In addition to simplifying estimation, it retains the efficiency benefits of semiparametric DRL while achieving the robustness guarantees of nonparametric DRL.
Our key insight is that the $Q$-function $q_{0}$ provides a sufficient one-dimensional summary for the Bellman equation, which enables us to reduce the conditional moment condition to
$q_{0}(A_{0},S_{0})-\gamma\,E_{0}[V^{\pi}(q_{0})(S_{1})\mid q_{0}(A_{0},S_{0})]=E_{0}[Y_{0}\mid q_{0}(A_{0},S_{0})],$
so that the equation conditions on the scalar $q_{0}(A_{0},S_{0})$ rather than the full tuple $(A_{0},S_{0})$. This identity follows from the law of total expectation, since
$q_{0}(A_{0},S_{0})=E_{0}\left[E_{0}\left[Y_{0}+\gamma V^{\pi}(q_{0})(S_{1})\mid A_{0},S_{0}\right]\mid q_{0}(A_{0},S_{0})\right]=E_{0}\left[Y_{0}+\gamma V^{\pi}(q_{0})(S_{1})\mid q_{0}(A_{0},S_{0})\right].$
This dimension reduction naturally motivates inference for an oracle projection parameter defined in terms of $q_{0}$. Specifically, if $q_{0}$ were known a priori, we could target the parameter $\Psi_{q_{0}}\colon P\mapsto E_{P}[m(S_{0},A_{0},q_{P,q_{0}})]$, where, for each $q\in L^{\infty}(\lambda)$, we define the Bellman projection
@@ 5.1 Leveraging the Q-function for dimension reduction
$q_{P,q}:=\operatorname*{argmin}_{f\circ q;\ f\colon\mathbb{R}\rightarrow\mathbb{R}}E_{P}\left[\left\{Y_{0}-\mathcal{T}_{P,q}(f\circ q)(A_{0},S_{0})\right\}^{2}\right],$
[…]
We now introduce notation used in the theorem. For a feature map $\phi:\mathcal{A}\times\mathcal{S}\rightarrow\mathbb{R}^{m}$, define the Riesz representer $\alpha_{0,\phi}:=\operatorname*{argmin}_{f\in\overline{H}_{P_{0},\phi}}E_{0}\left[\left\{\mathcal{T}_{0,\phi}(f\circ\phi)(A_{0},S_{0})\right\}^{2}-2m(S_{0},A_{0},f\circ\phi)\right]$, where $\overline{H}_{P_{0},\phi}$ is the closure of the model class $H_{\phi}:=\{f\circ\phi\text{ with }f:\mathbb{R}^{m}\rightarrow\mathbb{R}\}\cap L^{\infty}(\lambda)$ with respect to $\|\cdot\|_{P_{0}}$. Let $\widetilde{\alpha}_{0,q_{n}}:=\operatorname*{argmin}_{f\in\overline{H}_{P_{0},q_{n}}}\left\|\mathcal{T}_{0,(q_{n},q_{0})}(\alpha_{0,(q_{n},q_{0})})-\mathcal{T}_{0,q_{n}}(f\circ q_{n})\right\|_{P_{0}},$ where $\mathcal{T}_{0,q_{n}}(\widetilde{\alpha}_{0,q_{n}})$ is the projection of $\mathcal{T}_{0,(q_{n},q_{0})}(\alpha_{0,(q_{n},q_{0})})$ onto the range $\mathcal{T}_{0,q_{n}}(\overline{H}_{q_{n}})$.
@@
    (Continuity and invertibility) For $\phi=(q_{0}),$ $(q_{n})$, and $(q_{0},q_{n})$, the following hold:
    1. (a)
      There exists $C<\infty$ such that $|E_{0}[m(S_{0},A_{0},q)]|\leq C\|q\|_{P_{0}}$ for all $q\in H_{\phi}$.
    2. (b)
      $\mathcal{T}_{0,\phi}:\overline{H}_{P_{0},\phi}\rightarrow\overline{H}_{P_{0},\phi}$ is continuous and invertible.
In the following condition, we define the random variable $D_{n,0}:=\mathcal{T}_{0,(q_{n},q_{0})}\left(\alpha_{0,(q_{n},q_{0})}\right)(A_{0},S_{0})$.
@@
    (Lipschitz continuity) There exists a constant $L\in(0,\infty)$ such that, for all sufficiently large $n$, the bivariate function $(t_{1},t_{2})\mapsto E_{0}\left[D_{n,0}\,\middle|\,q_{n}(A_{0},S_{0})=t_{1},\ q_{0}(A_{0},S_{0})=t_{2},\ \mathcal{D}_{n}\right]$ is almost surely Lipschitz continuous with Lipschitz constant $L$.

### v4

@@ 3.2 Model-robust extension and asymptotic theory
Interestingly, restricting only the reward function or only the transition kernel does not relax the intertemporal overlap requirement. In both cases, the EIF—and therefore the efficiency bound—continues to depend critically on the state–occupancy ratio $d_{0}$ in (2).
Theorem 3 (EIF with known reward or kernel).
Let $\Psi(P):=E_{0}\!\left[V^{\pi}(q_{P})(S_{0})\right]$ and assume the law of $S_{0}$ is known. The EIF of $\Psi$ when either the reward or kernel is known is:
  1. (i)
    ($r_{0}$ is known) $\varphi_{0,1}(S_{1},A_{0},S_{0})\;=\;d_{0}(A_{0},S_{0})\,\Big\{r_{0}(A_{0},S_{0})+\gamma V^{\pi}(q_{0})(S_{1})-q_{0}(A_{0},S_{0})\Big\},$
  2. (ii)
    ($P_{0}(S_{1}\mid A_{0},S_{0})$ is known) $\varphi_{0,2}(Y_{0},A_{0},S_{0})\;=\;d_{0}(A_{0},S_{0})\,\Big\{Y_{0}-r_{0}(A_{0},S_{0})\Big\}.$
3.4 Sieve estimation and model approximation error
Sieve methods are widely used in econometrics and statistics because they combine the simplicity and finite-sample stability of parametric models with the flexibility of nonparametric function approximation (shen1997methods; newey1997convergence; spnpsieve). Rather than fixing a single working model $H$, one may adopt a sieve strategy, in which model selection for the $Q$-function is performed over a sequence of linear spaces $H_{1}\subset H_{2}\subset\cdots\subset H_{k}\subset\cdots\subset H_{\infty}:=H$ of increasing complexity. For example, $H_{k}$ could be a finite-dimensional series space spanned by trinometric polynomial (JacksonOnAB), spline (Gordon1974BSPLINECA), or wavelet bases (Antoniadis_2007), or it could be infinite-dimensional, such as a space of additively decomposable functions in $k$ variables (stone1985additive; hastie1987generalized; zhang2022regression). Given $H_{n}:=H_{k(n)}$ for a potentially data-dependent sequence $k(n)$, a sieve DRL estimator of $\psi_{0}$ is constructed in the same manner as in Section 3.1, yielding $\psi_{n,H_{n}}$. The next theorem shows that the associated misspecification bias $\Psi_{H_{n}}(P_{0})-\Psi(P_{0})$ is second order in the sieve approximation error, extending related bounds from static settings (van2023adaptive; van2025automatic). In Appendix F, we derive the asymptotic theory for sieve DRL estimators.
Theorem 4 (Model approximation error).
Suppose that $q_{0}\in H_{0}$ for some submodel $H_{0}\subseteq H$ (possibly depending on $P_{0}$), and that Condition C1 holds for both $H$ and $H_{0}$. Then
$\Psi_{H_{n}}(P_{0})-\Psi(P_{0})=-\left\langle\mathcal{T}_{0}(\alpha_{0,H_{n}})-\mathcal{T}_{0}(\alpha_{0,H_{n,0}}),\mathcal{T}_{0}(q_{0,H_{n}})-\mathcal{T}_{0}(q_{0})\right\rangle_{P_{0}},$
where $H_{n,0}:=H_{n}\oplus H_{0}$ denotes the direct-sum linear model.
For sieve DRL, when $k(n)\uparrow\infty$, the theorem may be applied with $H_{0}:=H$, in which case $\alpha_{0,H_{n,0}}=\alpha_{0,H}$. If $k(n)$ instead converges to a fixed finite value $k_{0}$, one may take $H_{0}:=H_{k_{0}}$. Applying the result with $H_{n}=H$ further implies that standard DRL estimators incur only second-order bias under mild misspecification of $H$.
3.5 Estimation of nuisance functions
We now describe how the $Q$-function $q_{0,H}$ and the Riesz representer $\alpha_{0,H}$ can be estimated using flexible machine learning methods. A standard approach for estimating $q_{0,H}$ in reinforcement learning is fitted $Q$-iteration (FQI) (munos2008finite), an iterative regression algorithm for solving the Bellman equation (1). This equation implies that $q_{0,H}(A_{0},S_{0})$ can be recovered by regressing the pseudo-outcome $Y_{0}+\gamma V^{\pi}(q_{0,H})(S_{1})$ on $(A_{0},S_{0})$. FQI initializes $q_{n,H}^{(0)}:=0$ and repeatedly regresses the updated Bellman target $Y_{0}+\gamma V^{\pi}(q_{n,H}^{(k)})(S_{1})$ on $(A_{0},S_{0})$ over the model class $H$ to obtain $q_{n,H}^{(k+1)}$. Iteration stops when successive updates change little in $\ell^{2}$-norm or when the cross-validated risk stabilizes. The finite-iteration error decreases at rate $\gamma^{k}$, so only $O(\log n)$ iterations are typically required in practice (munos2008finite; agarwal2019reinforcement; fan2020theoretical; hu2025fast). When implemented with gradient descent or boosting, the Bellman target may be updated at each gradient step, often accelerating convergence (baird1999reinforcement). A full description appears in Algorithm 2 in Appendix B.
Estimation of the Riesz representer $\alpha_{0,H}$ can be formulated as a convex–concave min–max problem analogous to those used for occupancy ratio estimation (liu2018breaking; nachum2019dualdice; uehara2020minimax; dikkala2020minimax; kallus2020double). The objective in (6) can be rewritten as $\alpha_{0,H}=\operatorname*{argmin}_{\alpha\in H}\max_{f\in L^{\infty}(\lambda)}L_{0}(\alpha,f)$, where the objective function is
$\begin{aligned} L_{0}(\alpha,f)=E_{0}\Big[\{\alpha(A_{0},S_{0})\}^{2}&-2\gamma\,\alpha(A_{0},S_{0})\,V^{\pi}(\alpha)(S_{1})-2\,m(S_{0},A_{0},\alpha)\\ &-\frac{\gamma^{2}}{2}\big\{\{f(A_{0},S_{0})\}^{2}-2\,V^{\pi}(\alpha)(S_{1})\,f(A_{0},S_{0})\big\}\Big].\end{aligned}$ (8)
Given an estimator $\alpha_{n,H}$ of $\alpha_{0,H}$, the remaining nuisance component in $\mathcal{T}_{0}(\alpha_{0,H})$ may be obtained by regressing $V^{\pi}(\alpha_{n,H})(S_{1})$ on $(A_{0},S_{0})$. Min–max empirical risk minimization is studied in dikkala2020minimax and bennett2023minimax. If the inner maximization is taken over a linear model or a reproducing kernel Hilbert space (RKHS), the maximizer can be computed efficiently in closed form using methods such as kernel ridge regression (kallus2020double; olivas2025source).
4 Superefficient inference via Bellman calibration
4.1 Proposed estimator
The previous section showed that imposing semiparametric structure on the $Q$-function can relax intertemporal overlap requirements and improve efficiency. However, if this structure is misspecified, such estimators may suffer from bias. In contrast, fully nonparametric DRL estimators remain valid but can exhibit high variance—or even lose asymptotic normality—when overlap is limited. To address this, we propose nonparametric superefficient estimators that use the $Q$-function as a one-dimensional adjustment variable in the Bellman equation. This reduction replaces the usual high-dimensional overlap condition with a single-dimensional one. As a result, the estimators are superefficient: their limiting variances lie strictly below the nonparametric efficiency bound (le1953some; van2000asymptotic; benkeser2020nonparametric).
Our key insight is that the $Q$-function $q_{0}$ satisfies a dimension-reduced Bellman equation:
$q_{0}(A_{0},S_{0})=E_{0}\!\left[Y_{0}+\gamma V^{\pi}(q_{0})(S_{1})\mid q_{0}(A_{0},S_{0})\right],$ (9)
which conditions only on the scalar value $q_{0}(A_{0},S_{0})$ rather than on the full state–action pair $(A_{0},S_{0})$. This identity follows by taking the conditional expectation given $q_{0}(A_{0},S_{0})$ of both sides of the Bellman equation in (1), and then applying the law of total expectation. Thus, if $q_{0}$ were known, we could apply DRL using $\{q_{0}(a,s):(a,s)\in\mathcal{A}\times\mathcal{S}\}$ as a sufficient one-dimensional summary of the state–action space. The resulting estimator would be superefficient in the fully nonparametric model—because it is efficient in the corresponding dimension-reduced model—with a limiting variance that depends only on intertemporal overlap conditional on the value of the $Q$-function, rather than on overlap across the full state–action space. Since $q_{0}$ is unknown, we instead apply this approach with an estimator $q_{n}^{*}$ of $q_{0}$. At first glance, this procedure may appear invalid, as it seems highly sensitive to errors in $q_{n}^{*}$. However, we show that this is not the case: the error incurred by substituting $q_{n}^{*}$ for $q_{0}$ enters only at second order.
Our proposed superefficient estimator of $\Psi(P_{0})$ is the calibrated plug-in estimator
$\psi_{n}^{*}:=\frac{1}{n}\sum_{i=1}^{n}m(S_{0,i},A_{0,i},q_{n}^{*}),$
where $q_{n}^{*}$ is a Bellman-calibrated estimator of the $Q$-function $q_{0}$, constructed to satisfy the empirical Bellman equation $q_{n}^{*}(a,s)=E_{P_{n}}\!\left[\,Y_{0}+\gamma V^{\pi}(q_{n}^{*})(S_{1})\,\middle|\,q_{n}^{*}(A_{0},S_{0})=q_{n}^{*}(a,s)\right],$ which can be written explicitly as
$q_{n}^{*}(a,s)=\frac{\sum_{i=1}^{n}1\{q_{n}^{*}(A_{0,i},S_{0,i})=q_{n}^{*}(a,s)\}\,\{Y_{0,i}+\gamma V^{\pi}(q_{n}^{*})(S_{1,i})\}}{\sum_{i=1}^{n}1\{q_{n}^{*}(A_{0,i},S_{0,i})=q_{n}^{*}(a,s)\}}.$ (10)
As we show, calibration performs implicit debiasing by ensuring that $q_{n}^{*}$ satisfies (9) empirically. Given an initial estimator $q_{n}$, a Bellman-calibrated estimator $q_{n}^{*}$ can be constructed post hoc using fitted $Q$-calibration and isotonic regression, as described in Algorithm 1.
Before presenting the calibration algorithm, we briefly explain why enforcing (10) debiases the plug-in estimator. The key observation is that Bellman calibration makes $\psi_{n}^{*}$ equivalent to an oracle DRL estimator in the dimension-reduced model induced by $q_{n}^{*}$. In particular, (10) is equivalent to requiring orthogonality of the temporal-difference errors: for every transformation $f:\mathbb{R}\to\mathbb{R}$,
$\frac{1}{n}\sum_{i=1}^{n}f\!\left(q_{n}^{*}(A_{0,i},S_{0,i})\right)\Bigl\{Y_{0,i}+\gamma V^{\pi}(q_{n}^{*})(S_{1,i})-q_{n}^{*}(A_{0,i},S_{0,i})\Bigr\}=0.$ (11)
The Riesz representer under the dimension-reduced model is of the form $f\circ q_{n}^{*}$ for some choice of $f$. Consequently, (11) guarantees that the bias-correction term in the corresponding DRL estimator vanishes when evaluated with $q_{n}^{*}$ and the true Riesz representer. In principle, one could instead estimate the Riesz representer directly and construct a superefficient DRL estimator by applying an explicit bias correction. However, although this yields an estimator that is asymptotically equivalent to ours, it requires consistent estimation of the representer—typically via a minimax optimization procedure—whereas calibration removes the need to estimate it at all.
4.2 Fitted $Q$-calibration
Our proposed algorithm, fitted $Q$-calibration with isotonic regression, is given in Algorithm 1, where $\mathcal{F}_{\text{iso}}$ denotes the class of monotone non-decreasing (isotonic) functions. Python Code is provided in Appendix E. The method combines isotonic calibration (zadrozny2002transforming), originally developed for classification, with fitted $Q$-iteration (munos2008finite). Operationally, it is equivalent to running fitted $Q$-iteration on the dimension-reduced data $\{q_{n}(A_{0,i},S_{0,i}),\,Y_{0,i}\}$, using monotone functions as the regression class. Its convergence properties therefore follow directly from standard results on fitted $Q$-iteration and isotonic regression. In particular, if $q_{n}$ is trained on data independent of the calibration sample, Section 5 of hu2025fast shows (under mild conditions) that
$\|q_{n}^{*(K)}-q_{0}\|_{P_{0}}\;\lesssim\;\|q_{n}-q_{0}\|_{P_{0}}+O_{p}(n^{-1/3})+O(\gamma^{K}),$
where $\|q_{n}-q_{0}\|_{P_{0}}$ upper bounds the misspecification error of the monotone class, $O_{p}(n^{-1/3})$ is the statistical error from estimating a monotone function, and $O(\gamma^{K})$ is the finite-iteration error. Consequently, calibration does not degrade the performance of $q_{n}$ beyond a typically negligible $O_{p}(n^{-1/3})$ term, and it often improves performance by learning an optimal monotone transformation. The number of iterations $K$ may be selected using a stopping rule such as $\|q_{n}^{*(k+1)}-q_{n}^{*(k)}\|_{P_{n}}<\varepsilon$. Because each isotonic regression update is computationally inexpensive, a tolerance on the order of $\varepsilon=1/n$ works well in practice.
Algorithm 1 Fitted $Q$-Calibration with isotonic regression
0:  Initial estimator $q_{n}$, max iteration $K$
1:  Initialize $q_{n}^{*(0)}:=q_{n}$
2:  for $k=0,1,2,\dots,K-1$ do
3:   $f_{n}^{(k+1)}:=\displaystyle\operatorname*{argmin}_{f\in\mathcal{F}_{\text{iso}}}\sum_{i=1}^{n}\left\{Y_{0,i}+\gamma V^{\pi}(q_{n}^{*(k)})(S_{1,i})-f(q_{n}(A_{0,i},S_{0,i}))\right\}^{2}$
4:   $q_{n}^{*(k+1)}:=f_{n}^{(k+1)}\circ q_{n}$
@@ 4.2 Fitted $Q$-calibration
6:  return $q_{n}^{*(K)}$
Algorithm 1 enforces (10) by exploiting the fact that isotonic regression is equivalent to histogram regression on data-dependent bins of the initial $Q$-values (barlow1972isotonic; van2023causal). A single isotonic regression fits a step function over a partition of the one-dimensional set $\{q_{n}(a,s):(a,s)\in\mathcal{A}\times\mathcal{S}\}$, with the partition chosen to satisfy monotonicity. For $K=\infty$, the calibrated estimate $q_{n}^{*}(a,s)$ equals the empirical average of the Bellman target $Y_{0,i}+\gamma V^{\pi}(q_{n}^{*})(S_{1,i})$ over all observations whose initial values $q_{n}(A_{0,i},S_{0,i})$ fall in the same bin as $q_{n}(a,s)$. However, these are precisely the observations whose calibrated values $q_{n}^{*}(A_{0,i},S_{0,i})$ equal $q_{n}^{*}(a,s)$, ensuring (10). An alternative is fixed-bin histogram calibration (zadrozny2001obtaining; gupta2021distribution), but it requires manual bin selection, whereas isotonic regression performs binning automatically (van2025generalized).
To mitigate overfitting, the initial $Q$-function estimator $q_{n}$ in Algorithm 1 should be trained on data that is independent of the data used for calibration. This can be achieved via sample splitting: one part of the data is used to estimate $q_{n}$, and the remaining part is used to calibrate $q_{n}$ and compute the plug-in estimator of the linear functional. To improve data efficiency, we recommend cross-fitting $q_{n}$ and then calibrating the pooled out-of-fold predictions using all available data, as in van2024stabilized; vanderlaan2024automatic. A cross-fitted version of Algorithm 1 is given in Appendix C.
4.3 Asymptotic theory
We show that the calibrated plug-in estimator is efficient for an oracle parameter $\Psi_{q_{0}}$, from which superefficiency for $\Psi$ follows. Following the framework in Section 3.2, for any $q\in L^{\infty}(\lambda)$, we define the projection parameter
@@ 4.3 Asymptotic theory
$\Psi_{q}(P):=E_{P}\!\left[m(S_{0},A_{0},q_{P,q})\right],\qquad q_{P,q}:=\operatorname*{argmin}_{f\circ q:\,f\in L^{\infty}(\mathbb{R})}E_{P}\!\left[\{Y_{0}-\mathcal{T}_{P,q}(f\circ q)(A_{0},S_{0})\}^{2}\right],$
where $\mathcal{T}_{P,q}(f\circ q)(a,s):=f(q(a,s))-\gamma E_{P}[V^{\pi}(f\circ q)(S_{1})\mid q(A_{0},S_{0})=q(a,s)],$ and $L^{\infty}(\mathbb{R})$ denotes the class of bounded real-valued transformations. Although $\Psi_{q_{0}}$ identifies the same estimand $\psi_{0}$ as $\Psi$, since $\Psi_{q_{0}}(P_{0})=\Psi(P_{0})$, its efficiency bound is typically smaller and less sensitive to violations of intertemporal overlap. The following example illustrates this phenomenon in the case of the policy value.
Example 4 (Efficiency bound under dimension reduction).
For the map $m(a,s,q):=V^{b_{0}}(q)(s)$ with $\pi:=b_{0}$, the $P_{0}$-EIF of $\Psi$ is $(s,a,y,s^{\prime})\mapsto d_{0}(a,s)\{y+\gamma V^{b_{0}}(q_{0})(s^{\prime})-q_{0}(a,s)\}+V^{b_{0}}(q_{0})(s)-\psi_{0},$ where $d_{0}(a,s):=\sum_{t=0}^{\infty}\gamma^{t}\,\frac{d\mathbb{P}_{0}((A_{t},S_{t})=(a,s))}{d\mathbb{P}_{0}((A_{0},S_{0})=(a,s))}$. The $P_{0}$-EIF of $\Psi_{q_{0}}$ is obtained by replacing $d_{0}$ with the dimension-reduced occupancy ratio
$d_{0,q_{0}}(a,s):=\sum_{t=0}^{\infty}\gamma^{t}\,\frac{d\mathbb{P}_{0}(q_{0}(A_{t},S_{t})=q_{0}(a,s))}{d\mathbb{P}_{0}(q_{0}(A_{0},S_{0})=q_{0}(a,s))},$
since $d_{0,q_{0}}(A_{0},S_{0})=E_{0}[d_{0}(A_{0},S_{0})\mid q_{0}(A_{0},S_{0})]$. Thus, the efficiency bound of $\Psi_{q_{0}}$ depends only on intertemporal overlap of the one-dimensional process $\{q_{0}(A_{t},S_{t})\}_{t\in\mathbb{N}}$, rather than the full state–action process.
Although $\Psi_{q_{0}}$ depends on the unknown $q_{0}$, it can be approximated by replacing $q_{0}$ with the estimator $q_{n}^{*}$. The theorem below shows that $\Psi_{q_{n}^{*}}$ agrees with $\Psi_{q_{0}}$ up to second-order error under suitable conditions. The result does not rely on calibration: the theorem remains valid if $q_{n}^{*}$ is replaced by any estimator $q_{n}$. To state the result, we introduce some notation. For a feature map $\phi:\mathcal{A}\times\mathcal{S}\rightarrow\mathbb{R}^{m}$, define the induced model class $H_{\phi}:=\{f\circ\phi:f:\mathbb{R}^{m}\rightarrow\mathbb{R}\}\cap L^{\infty}(\lambda)$. Let $\alpha_{0,\phi}\in\overline{H}_{P_{0},\phi}$ denote the associated Riesz representer, defined by $E_{0}[m(A_{0},S_{0},q)]=\langle\mathcal{T}_{0,\phi}(\alpha_{0,\phi}),\,\mathcal{T}_{0,\phi}(q)\rangle_{P_{0}}$ for all $q\in\overline{H}_{P_{0},\phi}$. For shorthand, write $d_{0,\phi}:=\mathcal{T}_{0,\phi}(\alpha_{0,\phi})$.
@@
    (Invertibility) For $\phi=(q_{0}),$ $(q_{n}^{*})$, and $(q_{0},q_{n}^{*})$, $\mathcal{T}_{0,\phi}:\overline{H}_{P_{0},\phi}\rightarrow\overline{H}_{P_{0},\phi}$ is continuous and invertible.
@@
    (Lipschitz continuity) There exists $L\in(0,\infty)$ such that, for all sufficiently large $n$, the bivariate function $(t_{1},t_{2})\mapsto E_{0}\left[d_{0,(q_{n}^{*},q_{0})}(A_{0},S_{0})\,\middle|\,q_{n}^{*}(A_{0},S_{0})=t_{1},\ q_{0}(A_{0},S_{0})=t_{2},\ \mathcal{D}_{n}\right]$ is almost surely Lipschitz continuous with constant $L$.